A farmer moves along the boundary of a square field of side 10m in 40 seconds. What will be the magnitude of displacement of the farmer at the end of 2 minutes 20 seconds from his initial position?
1
Expert's answer
2012-06-12T09:01:41-0400
As a farmer moves along the boundary of a square field of side 10m in 40 seconds, then his velocity is

V = 10/40 = 0.25 m/s.

So, the magnitude of displacement of the farmer at the end of 2 minutes 20 seconds (or 60·2+20 = 140 seconds) from his initial position will be

S = 140·0.25 = 35 m.